OpenStudy (anonymous):

Solve: |x - 1/5| = 3 x = –14/5 and x = 16/5 x = 14/5 and x = –16/5 x = –14/5 and x = –16/5 x = 14/5 and x = 16/5

OpenStudy (anonymous):

x-1/5=3 and x-1/5=-3 x=3+(1/5) and x-(1/5)=-3 x=16/5 and x=-3+(1/5) x=16/5 and x=-14/5

OpenStudy (anonymous):

as \[\left| x \right| =\] x,x>0 -x,x<0

OpenStudy (anonymous):

thanks